OpenStudy (anonymous):

If P(A or B) = 0.8, P(A) = 0.1 and P(B) = 0.9, determine P(A and B).

myininaya (myininaya):

P(A or B)=P(A)+P(B)-P(A and B) P(A and B)=P(A)+P(B)-P(A or B) P(A and B)=0.1+0.9-0.8=.1+.1=.2